OpenStudy (anonymous):

The company discovered that it costs $37 to produce 3 calculators, $213 to produce 7, $658 to produce 12. Using the quadratic function, find the cost of producing 8 calculators.

OpenStudy (radar):

Review the widget problem that amistre64 worked out for you. This is another one just like it except their producing calculators.

OpenStudy (anonymous):

The quadratic funtion is f(x)=ax^2+bx+c , f(3)=9a+3b+c=37 f(7)=49a+7b+c=213 f(12)=144a+12b+c=658 , after the solving triple equation , a=5 b=-6 c=10 , so 64*5-6*8+10=272

OpenStudy (anonymous):

After 1 second, the missile is 224 feet in the air; after 2 seconds, it is 416 feet in the air. using the quadratic function, find the height in feet, of the missile after 10 seconds in the air. missile was launched from the ground

OpenStudy (anonymous):

Use this , \[V0*t-1/2*g*t ^{2}\] have double equation with plugging to t with 1 and 2 solve the equation and find initial velocity and gravity so put these to equation.

OpenStudy (anonymous):

yeah i have no idea.